If M(-2,2) is the midpoint of segment PQ and the coordinates of P are (-9,7), find the coordinates of Q?
Answer by stanbon(75887)

If M(-2,2) is the midpoint of segment PQ and the coordinates of P are (-9,7), find the coordinates of Q?
-----------------------------------
Each coordinate of the midpoint is the average of the two corresponding
coordinates of the end points.
---
Let the missing endpoint be (x,y)
x-coordinate: (x+(-9)/2 = -2
x-9 = -4
x = 5
----------------
y-coordinate: (y+7)/2 = 2
y+7 = 4
y = -3
------------
Other endpoint: (5,-3)
----------------------------------------
Cheers,
Stan H.